Projection Admin 11/02
To do this evening/ was completed: We do not have the new schedule yet. I did, however, update the info on the cue sheets for this Thursday since Godzilla info was incorrect by a long shot. Please do not add a new title to a formula section unless you have accurate timing/received content and can confirm. I added Godzilla info to the cue sheets and updated SE shows this week. Scheduled shows on the LMS that were missing.
Presentation Issues:
Preshow Clips/Playlist: Built playlists for Being Heumann (which will need to be techd tomorrow. I was already in conversation with these folks today to confirm we received all assets, which we did, and asked about slide placement/playlist preferences. They confirmed which slide is for the intro/preshow and which is for the Q&A, which is super helpful. I asked if they'd like us to tech it sooner than Wednesday (which is when our current set of keys open up) and they said yes, so I asked for a window for tomorrow morning 10am - 2pm, just to make sure everything looks good. We should be receiving those keys tonight.
Playlists built for Bringing Out the Dead and Big Boys. Again, these will need a quick tech. Reached out to John to confirm some info. Got the memos today so we're good.
Trailer Updates:
Special Events/ AV set up: Setup mics and stands in T3 and ran through/adjusted the playlist for Dark Room shorts tonight. Checked in with Cristina. Jack is all set for tonight.
Trailer Updates:
Special Events/ AV set up: Setup mics and stands in T3 and ran through/adjusted the playlist for Dark Room shorts tonight. Checked in with Cristina. Jack is all set for tonight.
Took another look at It's A Wonderful Life and added a masking cue. All set for tonight.
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ingested Last piece of Music Driven preshow content.
Ingested the Great Muppet Caper to the LMS.
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ingested Last piece of Music Driven preshow content.
Ingested the Great Muppet Caper to the LMS.
Received new keys for YCSE, Godzilla, Clayface
Continued DCPing NoBudge shorts.
Outgoing/Shipping (content):
Print/video Inspection:
Asset tracking doc updated/checked?: Yes. Cleaned through today.
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Yes
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Outgoing/Shipping (content):
Print/video Inspection:
Asset tracking doc updated/checked?: Yes. Cleaned through today.
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Yes
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment